Corporation's stock had a required return of 11.50% last year, when the risk-free rate was 5.50% and the market risk premium was 4.75%. Now suppose there is a shift in investor risk aversion, and the market risk premium increases by 2%. The risk-free rate and corporation's beta remain unchanged. What is corporation's new required return?


14.03%
14.38%
14.74%
15.10%
15.48%
Answer #1

required return=risk free rate+Beta*market risk premium

Currently:

11.5=5.5+Beta*4.75

Beta=(11.5-5.5)/4.75

=1.263157895

New market risk premium=(4.75+2)=6.75%

Hence new required return=5.5+(1.263157895*6.75)

=14.03%(Approx).



answered by: ANURANJAN SARSAM